What volume is occupied by 0.25 grams of oxygen gas measured at 25.0 deg C

and 100.66 kPa? ( 5 marks)

  ( universal gas constant= 8.314 Kpa. L mol-1 K-1)





1
Expert's answer
2021-02-04T07:03:10-0500

p = 100.66 kPa

T = 25 + 273 = 298 K

M(O2) = 32 g/mol

m = 0.25 g

"n = \frac{m}{M} \\

n(O_2) = \frac{0.25}{32} = 0.0078 \;mol"

Ideal Gas Law

"pV = nRT \\

V = \frac{nRT}{p} \\

= \frac{0.0078 \times 8.314 \times 298}{100.66} \\

= 0.192 \;L"

Answer: 192 mL